A police jeep is chasing a culprit going on motorbike.The motorbike crosses a turning at a speed of 72km/hr.The jeep follows it at a speed of 90km/hr.crossing the turning ten seconds later than the bike .Assuming that they travel at constant speeds,how far from the turning will the jeep catch up with the bike?

Grade:11

1 Answers

Arun
6 years ago
VP = 90 km/h = 25 m/s. VC = 72 km/h = 20 m/s. In 10 sec culprit reaches at point B from A. Distance converted by culprit S = vt = 20 × 10 = 200 m. At time t = 10 sec the police jeep is 200 m behind the culprit Time = s/v = 200 / 5 = 40 s. (Relative velocity is considered). In 40 s the police jeep will move from A to a distance S, where S = vt = 25 × 40 = 1000 m = 1.0 km away. ∴ The jeep will catch up with the bike, 1 km far from the turning.